OpenStudy (anonymous):

Show that this subset of R^2 {(a,c),(b,d)} is linearly independent if and only iff ad-bc =/=0

terenzreignz (terenzreignz):

This takes time, hang on...

OpenStudy (anonymous):

alright np i have been struggling with this one for awhile

terenzreignz (terenzreignz):

I can prove one way, still working on the other :)

OpenStudy (anonymous):

ok i actually have a step by step key as well but the answers do not make sense to me

terenzreignz (terenzreignz):

Oh, so you already have the answers? >.> I'll work on it anyway, what a coincidence that I'm perusing through some of these at the moment ;)

OpenStudy (anonymous):

ya i know the steps they say to make but it seems to jump. it says to do a case wehre a=/=0 and do row reduction then you have to do it with a=0 ,c=0 and when c=/= 0

terenzreignz (terenzreignz):

Ok, I can prove it now, are you interested? :)

OpenStudy (anonymous):

yes please

terenzreignz (terenzreignz):

Ok, here we go :) {(a,c), (b,d)} is in R^2, so a,b,c,d are all in R (just to clarify) Suppose ad - bc = 0 Then there exist scalars that we can multiply to (a,c) and (b,d), these scalars are d(a, c) and -c(b, d), and taking the sum of these (ad, cd) + (-bc, -cd) = (ad - bc, cd - cd) = (0, 0) Meaning the set is linearly dependent. So if ad - bc = 0, then the set if linearly dependent, therefore If the set is linearly independent, ad - bc =/= 0 Get it? I'll stop here, I'll continue when you tell me to :)

terenzreignz (terenzreignz):

By the way, this has the condition that not both d and c are 0, but I'll show that later ;) Any questions so far? May I proceed?

OpenStudy (anonymous):

yes please

terenzreignz (terenzreignz):

Did you understand the first part, though?

OpenStudy (anonymous):

ya. but why did you use -c?

OpenStudy (anonymous):

oh never mind i see

terenzreignz (terenzreignz):

OK. I'll proceed then :)

terenzreignz (terenzreignz):

Suppose this time that the set is linearly dependent. Then one of them is a linear combination of the other elements. WLOG, suppose that (a, c) = k(b, d) for some scalar k. Then (a, c) = (kb, kd) a = kb c = kd Therefore a/b = k c/d = k a/b = c/d ad = bc ad - bc = 0 Therefore, if the set is linearly dependent, then ad - bc = 0 Thus, if ad - bc =/= 0 then the set is linearly independent QED (end of proof) Any questions? :)

OpenStudy (anonymous):

haha nope tyvm I really appreciate it

terenzreignz (terenzreignz):

No problem. I'll show that if d and c are both 0, then the set is linearly dependent: Suppose d = c = 0 Then the set is {(a,0), (b,0)} But it's linearly dependent, since there is a scalar, -b/a such that (-b/a)(a, 0) + (b, 0) (-b, 0) + (b, 0) = (0, 0) :)

OpenStudy (anonymous):

:)